# The Buffer Mechanism for Multi-Step Information Reasoning in Language Models

### v2

@@ Appendix A Experimental settings
In our experiments (Section 3), the vocabulary size is set to $d=201$, and the hidden space dimension is set to $d_{m}=400$ and $d_{q}=d_{k}=d_{v}=64$. We use 200,000 2-step reasoning sequences (with sequence length equal to 13). The learning rate is set to 2e-5 and linearly warms up to 1e-4 within 400 epochs and then decays to 1e-5 within 3600 epochs. The batch size is set to 100. We use the AdamW optimizer with default parameters as set in PyTorch 2.3.0. We employ a gradient clipping strategy with torch.nn.utils.clip_grad_norm_(model.parameters(), max_norm=1).
@@
The task shown above can be completed within 2 hours with a single NVIDIA A100 GPU. For other more complex examples, they can be finished within 24 hours.
Appendix B Proof of Lemma 1
@@
$\mathbb{E}[({\bm{W}}_{i}{\bm{W}}_{j}^{\mathsf{T}})_{s,t}]=\mathbb{E}[\sum_{k=1}^{d_{m}}({\bm{W}}_{i})_{sk}({\bm{W}}_{j})_{kt}]=\sum_{k=1}^{d_{m}}\mathbb{E}[({\bm{W}}_{i})_{sk}]\mathbb{E}[({\bm{W}}_{j})_{kt}]=0,$
$\mathrm{Var}\left[({\bm{W}}_{i}{\bm{W}}_{j}^{\mathsf{T}})_{s,t}\right]=\mathbb{E}\left[\left(\sum_{k=1}^{d_{m}}({\bm{W}}_{i})_{sk}({\bm{W}}_{j})_{kt}\right)^{2}\right]$
$=\sum_{k=1}^{d_{m}}\mathbb{E}\left[({\bm{W}}_{i})_{sk}^{2}\right]\mathbb{E}\left[({\bm{W}}_{j})_{kt}^{2}\right]=d_{m}\times(\frac{1}{d_{m}})^{2}=\frac{1}{d_{m}},$
@@
$\mathbb{E}_{\{{\bm{W}}_{j}\}}\left[({\bm{x}}{\bm{W}}_{i}^{\mathsf{T}})_{t}\right]$ $=\sum_{j=1}^{n}\mathbb{E}_{\{{\bm{Z}}^{(j,i)}\}_{j}}\left[({\bm{a}}_{j}{\bm{Z}}^{(j,i)})_{t}\right]$
$=\sum_{k=1}^{d_{m}}({\bm{a}}_{i})_{k}\mathbb{E}\left[({\bm{Z}}^{(i,i)})_{kt}\right]+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{d_{m}}({\bm{a}}_{j})_{k}\mathbb{E}\left[({\bm{Z}}^{(j,i)})_{kt}\right]$
$=\sum_{k=1}^{d_{m}}({\bm{a}}_{i})_{k}\mathbb{E}\left[({\bm{Z}}^{(i,i)})_{kt}\right]=({\bm{a}}_{i})_{t},$
@@
$=d_{m}\times\frac{1}{d_{m}}\times\frac{1}{d_{m}}+2\times d_{m}\times\frac{1}{d_{m}}\times\frac{n}{d_{m}}+d_{m}\times\frac{n}{d_{m}}\times\frac{n}{d_{m}}=\frac{(n+1)}{d_{m}}.$
@@
Appendix C Further Discussion on the Vertical Thinking Strategy
C.1 Information Fusion
In our symbolic dataset task, as mentioned in Section. 3, the first layer facilitates the fusion of token information at odd and even positions. We find that positional encoding plays a crucial role in the features of the first layer of attention. Fig. 9(a)(b) illustrates a comparison between the original attention mechanism and the positional attention mechanism calculated with eq. 18. As shown, there is minimal difference between the two approaches.
@@ C.1 Information Fusion
C.2 Detailed Matching Matrix
In Section. 3, for simplicity of analysis, we ignored the impact of the feedforward layer. Here, we define a detailed version of the matching matrix as follows:
@@ C.2 Detailed Matching Matrix
C.3 Weight Construction Method for Multi-Step Reasoning Networks
In Section. 3, we mentioned that by setting the weights in the following manner, we can enable an $L$-layer Transformer model to possess $(L-1)$-step reasoning capability.
@@ C.3 Weight Construction Method for Multi-Step Reasoning Networks
The specific construction method is as follows: To ensure that each buffer in the model has adequate robustness against interference, we set $d_{m}=d_{q}=d_{k}=d_{v}=10000$. The feedforward layers are assigned zero weights so that the residual connection dominates. Since all the weight matrices we use are untrained random matrices, the layer normalization will have no effect. Fig. 11 shows the multi-step reasoning ability of an 8-layer Transformer. We tested natural order, reverse order, random order sentences, and sentences with inserted irrelevant tokens (i.e., token [20]), and the model was able to output the correct answer [8] in all cases. Each sentence begins with token [20] to prevent $\mathcal{A}^{(l)}_{0,0}$ always equals to 1, which could affect the buffer.
@@
Appendix D Mechanism of Parallel Reasoning
In this section, we investigate a natural question arising from Section 3: Can the model still perform reasoning through the vertical thinking strategy when the number of reasoning steps required by the task exceeds the number of reasoning layers? Our experimental studies reveal that, in such cases, the model exhibits a parallel reasoning strategy. Based on these observations, we hypothesize that the upper bound of the model’s reasoning capability will grow exponentially with the depth of the model. We have conducted preliminary verification of this hypothesis using a simple program.
Mechanism of Parallel Reasoning. Fig. 12(a) illustrates the information flow of a 4-layer model completing 4-step reasoning. We observe that when the reasoning steps exceed or equal the number of model layers, the model performs reasoning parallelly in one layer. In the 2nd layer, different information undergoes two reasoning operations, enabling the subsequent three layers to achieve 4-step reasoning. Moreover, starting from the 2nd layer, the model’s reasoning approach transits from matching tokens’ values to matching tokens’ positions. Token value matching is only utilized in the 1st layer.
Figure 12: (a) Information flow of a 4-layer Transformer model completing 4-step reasoning. (b) Number of stored information tokens in the last position with respect to the number of iterations when propagating information based on the Info-Prop Rules. We randomly selected 1000 sentences for the simulation. The blue dots represent the simulation results. (c) The coefficients by which each information is multiplied.
Based on the mechanism of information propagation, at the $l$-th layer, if the information is transmitted through attention, it should be multiplied by a coefficient ${\bm{W}}^{vo(l)}$, whereas no such term is required if it is transmitted through residual connections. Consequently, in Fig. 12(c), we present the coefficients by which each information is multiplied. It is observed that for each positional information, the coefficients are distinct, and the same holds true for the value information. The variation in coefficients also indicates that the information is stored in different subspaces, thereby enabling parallel reasoning.
Upper Bound of Model’s Reasoning Ability. Through the study of the buffer mechanism in Transformer models, we can estimate the upper bound of the reasoning ability for a general L-layer Transformer network. First, we assume that the model satisfies the following assumption.
Assumption 1.
Assume that the hidden space dimension $d_{m}$ is sufficiently large, allowing different information to be stored in independent buffers with little interference. Moreover, information transmission is achieved solely through the same-token matching of the Attention module, without considering the influence of the feedforward module.
We regard $n$ tokens of a sequence as $n$ nodes, each storing position and value information. The Transformer model follows the Information Propagation (Info-Prop) Rules (denoting the node transmitting information as $i$ and the node receiving information as $j$, considering the existence of the attention mask, we require $i<j$):
  • Rule 1: Odd positions can pass information to subsequent even positions, i.e., node $i$ stores an odd number $a$ in its position information, and node $j$ stores $a+1$ in its position information.
  • Rule 2: The position information stored in node $i$ and node $j$ has common elements.
  • Rule 3: The value information stored in node $i$ and node $j$ has common elements.
If any of the above three rules are satisfied, we consider that the information from node $i$ can be transmitted to node $j$. When information transmission occurs, the position and value information stored in node $j$ will take the union of the corresponding information from node $i$ and node $j$.
Pseudocode 1 provides a detailed process of the Info-Prop Rules. Each Node[i] is a dictionary containing two attributes: pos (position) and val (value). Node[i][pos] and Node[i][val] are two sets. Initially, each set contains only its own information and will be updated according to the Info-Prop Rules. Fig. 12(b) shows the number of value information stored in the last token with respect to the number of iterations under our set of rules. It can be observed that the stored information exhibits an approximately exponential growth pattern with the number of iterations.
Algorithm 1 Information Propagation Rules
1:  Initialize an empty dictionary Node
2:  for each token i in the sequence do
3:     Node[i][pos] $\leftarrow$ i
4:     Node[i][val] $\leftarrow$ token i
5:  end for
6:  Node_new $\leftarrow$ Node
7:  Information propagation
8:  for each node i in Node do
9:     for each node j $>$ i in Node do
10:        rule1 $\leftarrow$ any even number $a\in$ Node[i][pos] satisfies $a+1\in$ Node[j][pos]
[…]
13:        if rule1 or rule2 or rule3 then
14:           Node_new[j][pos] $\leftarrow$ Node_new[j][pos] $\cup$ Node[i][pos]
15:           Node_new[j][val] $\leftarrow$ Node_new[j][val] $\cup$ Node[i][val]
16:        end if
17:     end for
18:  end for
A possible explanation for this experimental result is that in the $L$-th layer, the buffers of the Transformer can be represented as $\{\prod_{l\in J}{\bm{W}}^{vo(l)}|J\subset\{0,1,\cdots,L-1\}\}$. Therefore, the maximum amount of information (target information and position information) in the last layer is $2^{L+1}$. In practice, the hidden space dimensions $d_{m}$, $d_{q}$, $d_{k}$, and $d_{v}$ of large language models are far from meeting the requirements of Assumption 1. Consequently, the actual reasoning ability of large language models is constrained by the hidden space dimensions.
@@ Appendix E Details for RMBA Experiment
We replace ${\bm{W}}^{qk(l)}$ and ${\bm{W}}^{vo(l)}$ with ${\bm{W}}^{qk(l)}+\alpha^{(l)}{\bm{Z}}^{(l-1)}$ and ${\bm{W}}^{vo(l)}+\beta^{(l)}{\bm{Z}}^{(l)}$, respectively, where $\alpha^{(l)}$ and $\beta^{(l)}$ are learnable parameters, and $\{{\bm{Z}}^{(l)}\}_{l=1}^{L}$ is a set of random matrix following $N(0,1/d_{m})$. Therefore, 6 extra learnable parameters are added to this 3-layer single-head model in total. Fig. 13 and Fig. 14 show the loss and accuracy of the Transformer under different settings. Fig. 15 shows the changes of the learnable parameters $\alpha^{(l)}$ and $\beta^{(l)}$ during training.
@@
Figure 14: The average accuracy of different initial value of $\alpha^{(l)}$ and $\beta^{(l)}$.
(a) layer 0
(b) layer 1
(c) layer 2

### v3

@@ Appendix B Experimental settings
In our experiments (Section 3), the vocabulary size is set to $d=201$, and the hidden space dimension is set to $d_{m}=400$ and $d_{q}=d_{k}=d_{v}=64$. We use 200,000 2-step reasoning sequences (with sequence length equal to 13). The learning rate is set to 2e-5 and linearly warms up to 1e-4 within 400 epochs and then decays to 1e-5 within 3600 epochs. The batch size is set to 100. We use the AdamW optimizer with default parameters as set in PyTorch 2.3.0.
@@
The task shown above can be completed within 2 hours with a single NVIDIA A100 GPU. For other, more complex examples, they can be finished within 24 hours.
Appendix C Proof of Lemma 4.1
@@
$\mathrm{e}[({\bm{W}}_{i}{\bm{W}}_{j}^{\mathsf{T}})_{s,t}]=\mathrm{e}[\sum_{k=1}^{d_{m}}({\bm{W}}_{i})_{sk}({\bm{W}}_{j})_{kt}]=\sum_{k=1}^{d_{m}}\mathrm{e}[({\bm{W}}_{i})_{sk}]\mathrm{e}[({\bm{W}}_{j})_{kt}]=0,$
$\mathrm{Var}\left[({\bm{W}}_{i}{\bm{W}}_{j}^{\mathsf{T}})_{s,t}\right]=\mathrm{e}\left[\left(\sum_{k=1}^{d_{m}}({\bm{W}}_{i})_{sk}({\bm{W}}_{j})_{kt}\right)^{2}\right]$
$=\sum_{k=1}^{d_{m}}\mathrm{e}\left[({\bm{W}}_{i})_{sk}^{2}\right]\mathrm{e}\left[({\bm{W}}_{j})_{kt}^{2}\right]=d_{m}\times(\frac{1}{d_{m}})^{2}=\frac{1}{d_{m}},$
@@
$\mathrm{e}_{\{{\bm{W}}_{j}\}}\left[({\bm{x}}{\bm{W}}_{i}^{\mathsf{T}})_{t}\right]$ $=\sum_{j=1}^{n}\mathrm{e}_{\{{\bm{Z}}^{(j,i)}\}_{j}}\left[({\bm{a}}_{j}{\bm{Z}}^{(j,i)})_{t}\right]$
$=\sum_{k=1}^{d_{m}}({\bm{a}}_{i})_{k}\mathrm{e}\left[({\bm{Z}}^{(i,i)})_{kt}\right]+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sum_{k=1}^{d_{m}}({\bm{a}}_{j})_{k}\mathrm{e}\left[({\bm{Z}}^{(j,i)})_{kt}\right]$
$=\sum_{k=1}^{d_{m}}({\bm{a}}_{i})_{k}\mathrm{e}\left[({\bm{Z}}^{(i,i)})_{kt}\right]=({\bm{a}}_{i})_{t},$
@@
$=d_{m}\cdot\frac{1}{d_{m}}\cdot\frac{1}{d_{m}}+2\cdot d_{m}\cdot\frac{1}{d_{m}}\cdot\frac{n}{d_{m}}+d_{m}\cdot\frac{n}{d_{m}}\cdot\frac{n}{d_{m}}=\frac{(n+1)}{d_{m}}.$
@@
Appendix D Further Discussion on the Vertical Thinking Strategy
D.1 Information Fusion
In our symbolic dataset task, as mentioned in Section 3, the first layer facilitates the fusion of token information at odd and even positions. We find that positional encoding plays a crucial role in the features of the first layer of attention. Fig. 9(a)(b) illustrates a comparison between the original attention mechanism and the positional attention mechanism calculated with eq. 9. As shown, there is minimal difference between the two approaches.
@@ D.1 Information Fusion
D.2 Detailed Matching Matrix
In Section 3, for simplicity of analysis, we ignored the impact of the feedforward layer. Here, we define a detailed version of the matching matrix as follows:
@@ D.2 Detailed Matching Matrix
D.3 Independence of Buffers
To verify the independence between buffers, we computed and visualized the cosine similarity between row vectors of different buffers (${\bm{W}}^{vo(0)},{\bm{W}}^{vo(1)},{\bm{W}}^{vo(2)}$ and $\bm{I}$}). As shown in Fig. 11, apart from exhibiting a certain similarity within itself (so-called condense phenomenon(Luo et al., 2021)), each buffer remains nearly orthogonal to the others.
Figure 11: The cosine similarity between row vectors of different buffers (${\bm{W}}^{vo(0)},{\bm{W}}^{vo(1)},{\bm{W}}^{vo(2)}$ and $\bm{I}$}).
D.4 Weight Construction Method for Multi-Step Reasoning Networks
In Section 3, we mentioned that by setting the weights in the following manner, we can enable an $L$-layer Transformer model to possess (L1)(L-1)-step reasoning capability.
@@ D.4 Weight Construction Method for Multi-Step Reasoning Networks
The specific construction method is as follows: To ensure that each buffer in the model has adequate robustness against interference, we set $d_{m}=d_{q}=d_{k}=d_{v}=10000$. The feedforward layers are assigned zero weights so that the residual connection dominates. Since all the weight matrices we use are untrained random matrices, the layer normalization will have no effect. Fig. 12 shows the multi-step reasoning ability of an 8-layer Transformer. We tested natural order, reverse order, random order sentences, and sentences with inserted irrelevant tokens (i.e., token [20]), and the model was able to output the correct answer [8] in all cases. Each sentence begins with token [20] to prevent $\mathcal{A}^{(l)}_{0,0}$ from always equaling 1, which could affect the buffer.
@@ Appendix E Details for RMBA Experiment
We replace ${\bm{W}}^{qk(l)}$ and ${\bm{W}}^{vo(l)}$ with ${\bm{W}}^{qk(l)}+\alpha^{(l)}{\bm{Z}}^{(l-1)}$ and ${\bm{W}}^{vo(l)}+\beta^{(l)}{\bm{Z}}^{(l)}$, respectively, where $\alpha^{(l)}$ and $\beta^{(l)}$ are learnable parameters, and $\{{\bm{Z}}^{(l)}\}_{l=1}^{L}$ is a set of random matrix following $N(0,1/d_{m})$. Therefore, 6 extra learnable parameters are added to this 3-layer single-head model in total. Fig. 13 shows the loss of the Transformer under different settings. Fig. 14 shows the changes of the learnable parameters $\alpha^{(l)}$ and $\beta^{(l)}$ during training.
@@
We further tested the performance of the three algorithms under different hyperparameter configurations. We investigated the effects of weight decay, learning rate, and hidden dimension on training. For the RMBA and IMBA algorithms, we set $\alpha^{(l)}_{\text{ini}}$$\beta^{(l)}_{\text{ini}}$=0.01. As shown in Fig. 15, under various settings, the RMBA algorithm consistently facilitated the model’s ability to generalize.
Figure 15: A comparison of the training results for RMBA, IMBA, and the Baseline model under different hyperparameters is presented. We investigated learning rates ranging from 2e-5 to 1e-4 and weight decay values from 0.01 to 1. We considered four configurations of the hidden space dimension. For each hyperparameter setting, we conducted experiments using 3 different random seeds, totaling 720 experiments.
Figure 16: Examples from each dataset, along with their training curves.
Appendix F Details for Horizontal Thinking Strategy
We trained a 2-layer Transformer model and investigated its ability to perform multi-step reasoning with horizontal thinking. Specifically, we trained the Transformer model on 20,000 samples of length 13, each labeled with the result of a one-step reasoning process.
As shown in Fig. 17, when we fed the model’s output back into the model, it was able to generate the next step’s reasoning result, even though it had never been exposed to sentences longer than length 13 during training.
Figure 17: An illustration of the process of performing 3-step reasoning using a 2-layer Transformer model with CoT. The width of the connections in the diagram is based on the attention weights.
Appendix G Matching Score and Kernel Score for Real World LLMs
In this section, we calculate the matching score and kernel score of the large language model Phi-3(Abdin et al., 2024).
We focus on whether ${\bm{W}}^{vo}$ and ${\bm{W}}^{qk}$ function as the information buffer and the information extractor, respectively. To simplify our analysis, we temporarily disregard the effects of the feedforward layers. Following the method described in the main text, we compute $\text{Ker}^{(l_{1},l_{2})}={\bm{W}}^{qk(l_{1})}{\bm{W}}^{vo(l_{2}),\mathsf{T}}$ and observe whether it exhibits a dominant diagonal characteristic. For multi-head models, the above equation is modified as:
$\text{Ker}^{(l_{1},l_{2})}=\left(\sum_{h}{\bm{W}}^{q(l_{1},h)}{\bm{W}}^{k(l_{1},h),\mathsf{T}}\right){\bm{W}}^{vo(l_{2}),\mathsf{T}}.$ (14)
We define the Kernel Score (KS) as
$\text{KS}(\text{Ker}^{(l_{1},l_{2})})=\text{Trace}(\sigma(\text{Ker}^{(l_{1},l_{2})}))/d_{m},$ (15)
which measures the ability of layer $l_{1}$ in the model to extract information cached at layer $l_{2}$. As shown in Fig. 18(a), when $l_{1}\geq l_{2}$, the kernel score is nearly zero, which aligns with the logical sequence of information processing. When $l_{1}<l_{2}$, the kernel score decreases as $(l_{2}-l_{1})$ increases, indicating that the model tends to extract the most recently acquired information for further processing. In Fig. 18(b), we plot $\sigma(\text{Ker}^{(l_{1},l_{2})})$ and highlight the regions where the Kernel Score>0.3>0.3.
To further verify that the diagonal structure of $\text{Ker}^{(l_{1},l_{2})}$ arises from the alignment of model weights rather than the intrinsic diagonal structure of $\sum_{h}{\bm{W}}^{q(l_{1},h)}{\bm{W}}^{k(l_{1},h),\mathsf{T}}$ and ${\bm{W}}^{vo(l_{2}),\mathsf{T}}$, we conducted a small experiment, as illustrated in Fig. 18(d). In this experiment, we assume that both matrices $A$ and $B$ are noise-added identity matrices, where the noise scale is denoted by $\alpha$. We then compute the kernel scores of $A$, $B$, and their product $C=AB$. The results show that when $\max(\text{KS}(A),\text{KS}(B))<0.3$, the Kernel score of the product, $\text{KS}(C)$, is less than 0.025. However, as shown in Fig. 18(c), we observe that for many heads, even when $\max(\text{KS}(\sum_{h}{\bm{W}}^{q(h)}{\bm{W}}^{k(h),\mathsf{T}}),\text{KS}({\bm{W}}^{vo,\mathsf{T}}))<0.3$, the $\text{Ker}^{(l_{1},l_{2})}$ still has a large kernel score. This indicates that the alignment of model weights is the key factor driving the diagonal structure of $\text{Ker}^{(l_{1},l_{2})}$. In Fig. 18(f), we present an example where both $\sum_{h}{\bm{W}}^{q(l_{1},h)}{\bm{W}}^{k(l_{1},h),\mathsf{T}}$ and ${\bm{W}}^{vo(l_{2}),\mathsf{T}}$ appear relatively disordered individually, but their product exhibits a clear diagonal structure.
Another straightforward method is to directly set the diagonal elements of $\sum_{h}{\bm{W}}^{q(l_{1},h)}{\bm{W}}^{k(l_{1},h),\mathsf{T}}$ and ${\bm{W}}^{vo(l_{2}),\mathsf{T}}$ to zero, and then calculate the kernel score based on the resulting $\tilde{\text{Ker}}^{(l_{1},l_{2})}$. Fig. 18(e) illustrates this result, showing that it is approximately the same as that in Fig. 18(a).
Moreover, without loss of generality, we consider the case that includes LayerNorm(LN) and feedforward(FNN) layers. We compute the matching score for each head in each layer, with the specific calculation formula as follows:
${\bm{X}}^{vo}$ $=\text{LN}^{(l-1)}_{\text{attn}}({\bm{X}}){\bm{W}}^{v(l-1),\mathsf{T}}{\bm{W}}^{o(l-1),\mathsf{T}},$ (16)
${\bm{X}}^{vof}$ $={\bm{X}}^{vo}+\text{FNN}^{(l-1)}(\text{LN}^{(l-1)}_{\text{FNN}}({\bm{X}}^{vo})),$ (17)
${\bm{X}}^{vok(h)}$ $=\text{LN}^{(l)}_{\text{attn}}({\bm{X}}^{vof}){\bm{W}}^{k(l,h)},$ (18)
${\bm{X}}^{f}$ $=\text{LN}^{(l-1)}_{\text{attn}}({\bm{X}})+\text{FNN}^{(l-1)}(\text{LN}^{(l-1)}_{\text{FNN}}(\text{LN}^{(l-1)}_{\text{attn}}({\bm{X}}))),$ (19)
${\bm{X}}^{q(h)}$ $={\bm{X}}^{f}{\bm{W}}^{q(l,h)},$ (20)
$\text{matching matrix: }h^{(l,h)}({\bm{X}})$ $={\bm{X}}^{q(h)}{\bm{X}}^{vok(h),\mathsf{T}}.$ (21)
We visualized the matching score of each head in each layer (Fig. 18(b)) and found that the matching scores were highest in layers 5 to 20. This aligns with the conclusion mentioned in Dutta et al. (2024), namely that the reasoning layers of large language models generally appear in the middle portion.
To further validate that Phi-3 might employ a buffer mechanism, we computed the pairwise cosine similarity of the matrices $\{{\bm{W}}^{vo(l)}\}$ (each matrix is flattened as a long vector). The results indicate that these matrices are nearly orthogonal to each other, suggesting that they can be treated as independent buffers.
Figure 18: The calculation results of the kernel score and matching score for the Phi-3 model. (a) Visualization of $\sigma(\text{Ker}^{(l_{1},l_{2})})$. (b) Visualization of the matching score calculations for each head in each layer, indicating that reasoning is concentrated in the middle layers of the model. (c) Visualization of the kernel matrix between layers, where the subgraphs enclosed in red boxes correspond to $\text{KS}(\text{Ker})>0.3$ and the subgraphs enclosed in darkred boxes correspond to $\text{KS}(\text{Ker})>0.3$ but $\max(\text{KS}(\sum_{h}{\bm{W}}^{q(h)}{\bm{W}}^{k(h),\mathsf{T}}),\text{KS}({\bm{W}}^{vo,\mathsf{T}}))<0.3$. (d) The kernel score of a noise-added identity matrix(A and B) and the kernel score of the product of two noise-added identity matrices(C). It can be observed that for two unrelated matrices, when their individual kernel scores are less than 0.3, the kernel score of their product is less than 0.025. (e) The kernel score obtained after setting the diagonal elements of $\sum_{h}{\bm{W}}^{q(h)}{\bm{W}}^{k(h),\mathsf{T}}$ and ${\bm{W}}^{vo,\mathsf{T}}$ to zero. (f) Visualization of the $\sum_{h}{\bm{W}}^{q(14,h)}{\bm{W}}^{k(14,h),\mathsf{T}}$ and the ${\bm{W}}^{vo(l_{2}),\mathsf{T}}$ in the Phi-3 model, along with their inner product. Despite their weak diagonal structure individually, their inner product exhibits a clear diagonal structure. (g) Cosine similarity of flattened ${\bm{W}}^{vo(l_{1})}$ and ${\bm{W}}^{vo(l_{2})}$, $l_{1},l_{2}\in\{0,\cdots,31\}$.
Except for the phi-3 large model, the above definitions of the multi-head matching score and kernel score are equally applicable to other models that do not employ group query attention. The corresponding experimental results are shown in Fig. 19.
Figure 19: Weight alignment phenomenon for real-world LLMs.
Appendix H Causal Intervention
This section presents causal intervention experiments conducted to verify that the model uses a buffer mechanism when performing symbolic multi-step reasoning tasks. We assume readers are familiar with the content of Section 3. The causal intervention experiments were conducted using a 3-layer Transformer model trained as described in Section 5.
First, we identified critical tokens and logical circuits by observing output changes when masking specific attention or residual paths. Fig. 20(left) illustrates the logical circuit of the 3-layer Transformer performing the symbolic 2-step reasoning task. Subsequently, we conducted causal intervention experiments on the information stored in each token.
Figure 20: Logical circuit for 2-step reasoning (left) and illustration of causal intervention(right). We achieve the goal of intervening in information stored in a specific buffer by modifying the input sequence and enforcing the same attention pattern as before. The figure illustrates how the information [a] stored in the final token’s buffer ${\bm{W}}^{vo(0)}{\bm{W}}^{vo(1)}$ in layer 2 can be changed to [x].
Unlike prior works where causal intervention replaced all information within a token with alternative information(Feng and Steinhardt, 2023; Meng et al., 2022; Vig et al., 2020; Wang et al., 2024a), we refined the perturbation scope to target a specific buffer within a token. Specifically, we individually replaced the information in each buffer of critical tokens with alternative information to observe the model’s output. For example, as shown in Fig. 20(right), suppose we want to change the information [a] stored in the final token’s buffer ${\bm{W}}^{vo(0)}{\bm{W}}^{vo(1)}$ in layer 2 to [x]. This can be achieved by simply modifying the input sentence [a’][b’][b’][c’][a][b][b][c]...[a] to [a’][b’][b’][c’][x][b][b][c]...[a], and enforcing $\text{Attn}^{(1)}_{[13:]}$ same as before. Fig. 21 shows the intervention results for all buffers of all critical tokens.
The results reveal that, in layer $l$, the information stored in buffer $W^{vo(l-1)}$ of the final token is crucial. Modifying information in other buffers does not affect the model’s output. Combined with the observation in Appendix D that the cosine similarities between ${\bm{W}}^{vo(0)},{\bm{W}}^{vo(1)},{\bm{W}}^{vo(2)},{\bm{I}}$ are nearly zero, we can confidently assert that the model performs reasoning by leveraging different buffers. This experiment also rules out the possibility of the model only using an overwrite mechanism to perform reasoning.
Figure 21: Causal Intervention Experiment. We performed interventions on the information stored in every buffer of every critical token individually. Here, [x] represents a token that does not appear in the original input. For the final token, only modifying the buffer ${\bm{W}}^{vo(l-1)}$ in layer $l$ affects the final output. In this experiment, the tokens in the original sentence are selected from the range [20, 40], while token [x] traverses the range [40, 100]. $\text{Attn}^{(l)}_{[13:]}$ refers to the attention score corresponding to the last token at layer $l$. For the original input, we have $\text{Attn}^{(1)}_{[13,6]}=1$ and $\text{Attn}^{(2)}_{[13,8]}=1$. Instances labeled as “Random" indicate that the output varies erratically as [x] changes. In all other cases, the probability of the model output deviating from the value presented in the table is less than 1e-15.
Appendix I Interaction Results with Large Language Models
Figure 22: Detailed interaction results with large models. For each type of reasoning task, we tested each large model 25 times. The versions of Claude are Claude 3.5-haiku-20241022 and Claude 3.5-sonnet-20241022.